show that the electrical force between the electron and proton in the hydrogen atom is approximately 2.2e39 times stronger than the gravitational force between them.
Answer this question please
Thanks
1
Expert's answer
2019-12-05T10:22:51-0500

The absolute value of electric force between proton and electron in the hydrogen atom is "F_E = \frac{k e^2}{R^2}", where "k = 9 \cdot 10^9 N m^2 C^{-2}" is Coulomb constant, "e = 1.6 \cdot 10^{-19} C" is the charge of the electron, "R" is the radius of the hydrogen atom.

The absolute value of the gravitational force between proton and electron in the hydrogen atom is "F_G = \frac{G m_e m_p}{R^2}", where "G = 6.67 \cdot 10^{-11} m^3 kg^{-1}s^{-2}" is the gravitational constant, "m_e = 9.11 \cdot 10^{-31} kg" is the mass of the electron, "m_p = 1.67 \cdot 10^{-27} kg" is the mass of the proton.

The ratio of "F_E" to "F_G" is "\frac{k e^2}{R^2} \frac{R^2}{G m_e m_p} = \frac{k e^2}{G m_e m_p}".

Calculating with the values above, obtain "\frac{F_E}{F_G} \approx 2.2 \cdot 10^{39}".